Calculate the energy of a 25 Kg object sitting on a 50 meter edge.
1
Expert's answer
2020-11-16T07:45:34-0500

The potential energy of a body of mass "m=25kg" elevated on the height "h = 50m" above the ground is:


"W = mgh = 25kg\cdot 9.81m/s^2\cdot 50m = 12262.5\space J"

Here "g = 9.81m\/s^2" is the gravitational acceleration.


Answer. 12262.5 J.